GNU Readline Library

Brian Fox Free Software Foundation

Version 1.1 April 1991 This document describes the GNU Readline Library, a utility which aids in the consistency of user interface across discrete programs that need to provide a command line interface. Published by the Free Software Foundation 675 Massachusetts Avenue, Cambridge, MA 02139 USA Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc.

Appendix A Command Line Editing

This text describes GNU’s command line editing interface.

A.1 Introduction to Line Editing

The following paragraphs describe the notation we use to represent keystrokes.

The text ^{⟨}\underline{\overline{C-k}}^{⟩} is read as ‘Control-K’ and describes the character produced when the Control key is depressed and the ^{⟨}\underline{\overline{k}}^{⟩} key is struck.

The text ^{⟨}\underline{\overline{M-k}}^{⟩} is read as ‘Meta-K’ and describes the character produced when the meta key (if you have one) is depressed, and the ^{⟨}\underline{\overline{k}}^{⟩} key is struck. If you do not have a meta key, the identical keystroke can be generated by typing ^{⟨}\underline{\overline{ESC}}^{⟩} first, and then typing ^{⟨}\underline{\overline{k}}^{⟩}. Either process is known as metafying the ^{⟨}\underline{\overline{k}}^{⟩} key.

The text ^{⟨}\underline{\overline{M-C-k}}^{⟩} is read as ‘Meta-Control-k’ and describes the character produced by metafying ^{⟨}\underline{\overline{C-k}}^{⟩}.

In addition, several keys have their own names. Specifically, ^{⟨}\underline{\overline{DEL}}^{⟩}, ^{⟨}\underline{\overline{ESC}}^{⟩}, ^{⟨}\underline{\overline{LFD}}^{⟩}, ^{⟨}\underline{\overline{SPC}}^{⟩}, ⟨\underline{\overline{RET}}^{⟩}, and ^{⟨}\underline{\overline{TAB}}^{⟩} all stand for themselves when seen in this text, or in an init file (see Section A.3 [Readline Init File], page 3, for more info).

A.2 Readline Interaction

Often during an interactive session you type in a long line of text, only to notice that the first word on the line is misspelled. The Readline library gives you a set of commands for manipulating the text as you type it in, allowing you to just fix your typo, and not forcing you to retype the majority of the line. Using these editing commands, you move the cursor to the place that needs correction, and delete or insert the text of the corrections. Then, when you are satisfied with the line, you simply press ^{⟨}\underline{\overline{RET}}^{⟩}. You do not have to be at the end of the line to press ^{⟨}\underline{\overline{RET}}^{⟩}; the entire line is accepted regardless of the location of the cursor within the line.

A.2.1 Readline Bare Essentials

In order to enter characters into the line, simply type them. The typed character appears where the cursor was, and then the cursor moves one space to the right. If you mistype a character, you can use ^{⟨}\underline{\overline{DEL}}^{⟩} to back up, and delete the mistyped character.

Sometimes you may miss typing a character that you wanted to type, and not notice your error until you have typed several other characters. In that case, you can type ^{⟨}\underline{\overline{C-b}}^{⟩} to move the cursor to the left, and then correct your mistake. Aftwerwards, you can move the cursor to the right with ^{⟨}\underline{\overline{C-f}}^{⟩}.

When you add text in the middle of a line, you will notice that characters to the right of the cursor get ‘pushed over’ to make room for the text that you have inserted. Likewise, when you delete text behind the cursor, characters to the right of the cursor get ‘pulled back’ to fill in the blank space created by the removal of the text. A list of the basic bare essentials for editing the text of an input line follows.

⟨\underline{\overline{C-b}}^{⟩} Move back one character.

⟨\underline{\overline{C-f}}^{⟩} Move forward one character.

⟨\underline{\overline{DEL}}^{⟩} Delete the character to the left of the cursor.

⟨\underline{\overline{C-d}}^{⟩} Delete the character underneath the cursor.

Printing characters Insert itself into the line at the cursor.

⟨\underline{\overline{C-}} ^{⟩} Undo the last thing that you did. You can undo all the way back to an empty line.

A.2.2 Readline Movement Commands

The above table describes the most basic possible keystrokes that you need in order to do editing of the input line. For your convenience, many other commands have been added in addition to ^{⟨}\underline{\overline{C-b}}^{⟩}, ^{⟨}\underline{\overline{C-f}}^{⟩}, ^{⟨}\underline{\overline{C-d}}^{⟩}, and ^{⟨}\underline{\overline{DEL}}^{⟩}. Here are some commands for moving more rapidly about the line.

⟨\underline{\overline{C-a}}^{⟩} Move to the start of the line.

⟨\underline{\overline{C-e}}^{⟩} Move to the end of the line.

⟨\underline{\overline{M-f}}^{⟩} Move forward a word.

⟨\underline{\overline{M-b}}^{⟩} Move backward a word.

⟨\underline{\overline{C-l}}^{⟩} Clear the screen, reprinting the current line at the top.

Notice how^{⟨}\underline{\overline{C-f}}^{⟩} moves forward a character, while^{⟨}\underline{\overline{M-f}}^{⟩} moves forward a word. It is a loose convention that control keystrokes operate on characters while meta keystrokes operate on words.

A.2.3 Readline Killing Commands

Killing text means to delete the text from the line, but to save it away for later use, usually by yanking it back into the line. If the description for a command says that it ‘kills’ text, then you can be sure that you can get the text back in a different (or the same) place later.

Here is the list of commands for killing text.

⟨\underline{\overline{C-k}}^{⟩} Kill the text from the current cursor position to the end of the line.

⟨\underline{\overline{M-d}}^{⟩} Kill from the cursor to the end of the current word, or if between words, to the end of the next word.

⟨\underline{\overline{M-DEL}}^{⟩} Kill from the cursor to the start of the previous word, or if between words, to the start of the previous word.

⟨\underline{\overline{C-w}}^{⟩} Kill from the cursor to the previous whitespace. This is different than ^{⟨}\underline{\overline{M-DEL}}^{⟩} because the word boundaries differ.

And, here is how to yank the text back into the line.

⟨\underline{\overline{C-y}}^{⟩} Yank the most recently killed text back into the buffer at the cursor.

⟨\underline{\overline{M-y}}^{⟩} Rotate the kill-ring, and yank the new top. You can only do this if the prior command is ^{⟨}\underline{\overline{C-y}}^{⟩} or ^{⟨}\underline{\overline{M-y}}^{⟩}.

When you use a kill command, the text is saved in a kill-ring. Any number of consecutive kills save all of the killed text together, so that when you yank it back, you get it in one clean sweep. The kill ring is not line specific; the text that you killed on a previously typed line is available to be yanked back later, when you are typing another line.

A.2.4 Readline Arguments

You can pass numeric arguments to Readline commands. Sometimes the argument acts as a repeat count, other times it is the sign of the argument that is significant. If you pass a negative argument to a command which normally acts in a forward direction, that command will act in a backward direction. For example, to kill text back to the start of the line, you might type ^{⟨}\underline{\overline{M–}}^{⟩} ^{⟨}\underline{\overline{C-k}}^{⟩}.

The general way to pass numeric arguments to a command is to type meta digits before the command. If the first ‘digit’ you type is a minus sign (^{⟨}\underline{\overline{-}}^{⟩}), then the sign of the argument will be negative. Once you have typed one meta digit to get the argument started, you can type the remainder of the digits, and then the command. For example, to give the ^{⟨}\underline{\overline{C-d}}^{⟩} command an argument of 10, you could type ^{⟨}\underline{\overline{M-1 0 C-d}}^{⟩}.

A.3 Readline Init File

Although the Readline library comes with a set of gnu Emacs-like keybindings, it is possible that you would like to use a different set of keybindings. You can customize programs that use Readline by putting commands in an init file in your home directory. The name of this file is ‘~/.inputrc’.

When a program which uses the Readline library starts up, the ‘~/.inputrc’ file is read, and the keybindings are set.

In addition, the ^{⟨}\underline{\overline{C-x C-r}}^{⟩} command re-reads this init file, thus incorporating any changes that you might have made to it.

A.3.1 Readline Init Syntax

There are only four constructs allowed in the ‘~/.inputrc’ file:

Variable Settings You can change the state of a few variables in Readline. You do this by using the set command within the init file. Here is how you would specify that you wish to use vi line editing commands:

               set editing-mode vi

Right now, there are only a few variables which can be set; so few in fact, that we just iterate them here:

          editing-mode

The editing-mode variable controls which editing mode you are using. By default, gnu Readline starts up in Emacs editing mode, where the keystrokes are most similar to Emacs. This variable can either be set to emacs or vi.

horizontal-scroll-mode

This variable can either be set to On or Off. Setting it to On means that the text of the lines that you edit will scroll horizontally on a single screen line when they are larger than the width of the screen, instead of wrapping onto a new screen line. By default, this variable is set to Off.

mark-modified-lines

This variable when set to On, says to display an asterisk (‘*’) at the starts of history lines which have been modified. This variable is off by default.

prefer-visible-bell

If this variable is set to On it means to use a visible bell if one is available, rather than simply ringing the terminal bell. By default, the value is Off.

Key Bindings The syntax for controlling keybindings in the ‘~/.inputrc’ file is simple. First you have to know the name of the command that you want to change. The following pages contain tables of the command name, the default keybinding, and a short description of what the command does. Once you know the name of the command, simply place the name of the key you wish to bind the command to, a colon, and then the name of the command on a line in the ‘~/.inputrc’ file. The name of the key can be expressed in different ways, depending on which is most comfortable for you.

keyname: function-name or macro

keyname is the name of a key spelled out in English. For example:

     Control-u: universal-argument
     Meta-Rubout: backward-kill-word
     Control-o: ">&output"

In the above example, ^{⟨}\underline{\overline{C-u}}^{⟩} is bound to the function universalargument, and ^{⟨}\underline{\overline{C-o}}^{⟩} is bound to run the macro expressed on the right hand side (that is, to insert the text ‘>&output’ into the line).

"keyseq": function-name or macro keyseq differs from keyname above in that strings denoting an entire key sequence can be specified. Simply place the key sequence in double quotes. gnu Emacs style key escapes can be used, as in the following example:

     "\C-u": universal-argument
     "\C-x\C-r": re-read-init-file
     "\e[11~": "Function Key 1"

In the above example, ^{⟨}\underline{\overline{C-u}}^{⟩} is bound to the function universalargument (just as it was in the first example), ^{⟨}\underline{\overline{C-x C-r}}^{⟩} is bound to the function re-read-init-file, and ^{⟨}\underline{\overline{ESC [ 1 1 ~}}^{⟩} is bound to insert the text ‘Function Key 1’.

A.3.1.1 Commands For Moving

beginning-of-line (^{⟨}\underline{\overline{C-a}}^{⟩})

Move to the start of the current line.

end-of-line (^{⟨}\underline{\overline{C-e}}^{⟩})

Move to the end of the line.

forward-char (^{⟨}\underline{\overline{C-f}}^{⟩})

Move forward a character.

backward-char (^{⟨}\underline{\overline{C-b}}^{⟩})

Move back a character.

forward-word (^{⟨}\underline{\overline{M-f}}^{⟩})

Move forward to the end of the next word.

backward-word (^{⟨}\underline{\overline{M-b}}^{⟩})

Move back to the start of this, or the previous, word.

clear-screen (^{⟨}\underline{\overline{C-l}}^{⟩})

Clear the screen leaving the current line at the top of the screen.

A.3.1.2 Commands For Manipulating The History

accept-line (Newline, Return)

Accept the line regardless of where the cursor is. If this line is non-empty, add it to the history list. If this line was a history line, then restore the history line to its original state.

previous-history (^{⟨}\underline{\overline{C-p}}^{⟩})

Move ‘up’ through the history list.

next-history (^{⟨}\underline{\overline{C-n}}^{⟩})

Move ‘down’ through the history list.

beginning-of-history (^{⟨}\underline{\overline{M-<}}^{⟩})

Move to the first line in the history.

end-of-history (^{⟨}\underline{\overline{M->}}^{⟩})

Move to the end of the input history, i.e., the line you are entering.

reverse-search-history (^{⟨}\underline{\overline{C-r}}^{⟩})

Search backward starting at the current line and moving ‘up’ through the history as necessary. This is an incremental search.

forward-search-history (^{⟨}\underline{\overline{C-s}}^{⟩})

Search forward starting at the current line and moving ‘down’ through the the history as necessary.

A.3.1.3 Commands For Changing Text

delete-char (^{⟨}\underline{\overline{C-d}}^{⟩})

Delete the character under the cursor. If the cursor is at the beginning of the line, and there are no characters in the line, and the last character typed was not ^{⟨}\underline{\overline{C-d}}^{⟩}, then return EOF.

backward-delete-char (Rubout)

Delete the character behind the cursor. A numeric argument says to kill the characters instead of deleting them.

quoted-insert (^{⟨}\underline{\overline{C-q}}^{⟩}, ^{⟨}\underline{\overline{C-v}}^{⟩})

Add the next character that you type to the line verbatim. This is how to insert things like ^{⟨}\underline{\overline{C-q}}^{⟩} for example.

tab-insert (^{⟨}\underline{\overline{M-TAB}}^{⟩})

Insert a tab character.

self-insert (a, b, A, 1, !, ...)

Insert yourself.

transpose-chars (^{⟨}\underline{\overline{C-t}}^{⟩})

Drag the character before point forward over the character at point. Point moves forward as well. If point is at the end of the line, then transpose the two characters before point. Negative arguments don’t work.

transpose-words (^{⟨}\underline{\overline{M-t}}^{⟩})

Drag the word behind the cursor past the word in front of the cursor moving the cursor over that word as well.

upcase-word (^{⟨}\underline{\overline{M-u}}^{⟩})

Uppercase all letters in the current (or following) word. With a negative argument, do the previous word, but do not move point.

downcase-word (^{⟨}\underline{\overline{M-l}}^{⟩})

Lowercase all letters in the current (or following) word. With a negative argument, do the previous word, but do not move point.

capitalize-word (^{⟨}\underline{\overline{M-c}}^{⟩})

Uppercase the first letter in the current (or following) word. With a negative argument, do the previous word, but do not move point.

A.3.1.4 Killing And Yanking

kill-line (^{⟨}\underline{\overline{C-k}}^{⟩})

Kill the text from the current cursor position to the end of the line.

backward-kill-line ()

Kill backward to the beginning of the line. This is normally unbound.

kill-word (^{⟨}\underline{\overline{M-d}}^{⟩})

Kill from the cursor to the end of the current word, or if between words, to the end of the next word.

backward-kill-word (^{⟨}\underline{\overline{M-DEL}}^{⟩})

Kill the word behind the cursor.

unix-line-discard (^{⟨}\underline{\overline{C-u}}^{⟩})

Kill the whole line the way ^{⟨}\underline{\overline{C-u}}^{⟩} used to in Unix line input. The killed text is saved on the kill-ring.

unix-word-rubout (^{⟨}\underline{\overline{C-w}}^{⟩})

Kill the word the way ^{⟨}\underline{\overline{C-w}}^{⟩} used to in Unix line input. The killed text is saved on the kill-ring. This is different than backward-kill-word because the word boundaries differ.

yank (^{⟨}\underline{\overline{C-y}}^{⟩})

Yank the top of the kill ring into the buffer at point.

yank-pop (^{⟨}\underline{\overline{M-y}}^{⟩})

Rotate the kill-ring, and yank the new top. You can only do this if the prior command is yank or yank-pop.

A.3.1.5 Specifying Numeric Arguments

digit-argument (^{⟨}\underline{\overline{M-0}}^{⟩}, ^{⟨}\underline{\overline{M-1}}^{⟩}, ... ^{⟨}\underline{\overline{M–}}^{⟩})

Add this digit to the argument already accumulating, or start a new argument. ⟨\underline{\overline{M–}}^{⟩} starts a negative argument.

universal-argument ()

Do what ^{⟨}\underline{\overline{C-u}}^{⟩} does in gnu Emacs. By default, this is not bound.

A.3.1.6 Letting Readline Type For You

complete (TAB)

Attempt to do completion on the text before point. This is implementation defined. Generally, if you are typing a filename argument, you can do filename completion; if you are typing a command, you can do command completion, if you are typing in a symbol to GDB, you can do symbol name completion, if you are typing in a variable to Bash, you can do variable name completion.

possible-completions (M-?)

List the possible completions of the text before point.

A.3.1.7 Some Miscellaneous Commands

re-read-init-file (^{⟨}\underline{\overline{C-x}}^{⟩} ^{⟨}\underline{\overline{C-r}}^{⟩})

Read in the contents of your ‘~/.inputrc’ file, and incorporate any bindings found there.

abort (^{⟨}\underline{\overline{C-g}}^{⟩})

Stop running the current editing command.

prefix-meta (ESC)

Make the next character that you type be metafied. This is for people without a meta key. Typing ^{⟨}\underline{\overline{ESC f}}^{⟩} is equivalent to typing ^{⟨}\underline{\overline{M-f}}^{⟩}.

undo (^{⟨}\underline{\overline{C-}}^{˙ ⟩})

Incremental undo, separately remembered for each line.

revert-line (^{⟨}\underline{\overline{M-r}}^{⟩})

Undo all changes made to this line. This is like typing the ‘undo’ command enough times to get back to the beginning.

A.3.2 Readline vi Mode

While the Readline library does not have a full set of vi editing functions, it does contain enough to allow simple editing of the line.

In order to switch interactively between gnu Emacs and vi editing modes, use the command ^{⟨}\underline{\overline{M-C-j}}^{⟩} (toggle-editing-mode).

When you enter a line in vi mode, you are already placed in ‘insertion’ mode, as if you had typed an ‘i’. Pressing ^{⟨}\underline{\overline{ESC}}^{⟩} switches you into ‘edit’ mode, where you can edit the text of the line with the standard vi movement keys, move to previous history lines with ‘k’, and following lines with ‘j’, and so forth.

1 Programming with GNU Readline

This manual describes the interface between the GNU Readline Library and user programs. If you are a programmer, and you wish to include the features found in GNU Readline in your own programs, such as completion, line editing, and interactive history manipulation, this documentation is for you.

1.1 Default Behaviour

Many programs provide a command line interface, such as mail, ftp, and sh. For such programs, the default behaviour of Readline is sufficient. This section describes how to use Readline in the simplest way possible, perhaps to replace calls in your code to gets ().

The function readline prints a prompt and then reads and returns a single line of text from the user. The line which readline () returns is allocated with malloc (); you should free () the line when you are done with it. The declaration for readline in ANSI C is

     char *readline (char *prompt);

So, one might say

     char *line = readline ("Enter a line: ");

in order to read a line of text from the user.

The line which is returned has the final newline removed, so only the text of the line remains.

If readline encounters an EOF while reading the line, and the line is empty at that point, then (char *)NULL is returned. Otherwise, the line is ended just as if a newline was typed.

If you want the user to be able to get at the line later, (with ^{⟨}\underline{\overline{C-p}}^{⟩} for example), you must call add_history () to save the line away in a history list of such lines.

     add_history (line);

For full details on the GNU History Library, see the associated manual.

It is polite to avoid saving empty lines on the history list, since it is rare than someone has a burning need to reuse a blank line. Here is a function which usefully replaces the standard gets () library function:

     /* A static variable for holding the line. */
     static char *line_read = (char *)NULL;
     /* Read a string, and return a pointer to it.  Returns NULL on EOF. */
     char *
     do_gets ()
     {
       /* If the buffer has already been allocated, return the memory
          to the free pool. */
       if (line_read != (char *)NULL)
         {
           free (line_read);
           line_read = (char *)NULL;
         }

       /* Get a line from the user. */
       line_read = readline ("");
       /* If the line has any text in it, save it on the history. */
       if (line_read && *line_read)
         add_history (line_read);
       return (line_read);
     }

The above code gives the user the default behaviour of ^{⟨}\underline{\overline{TAB}}^{⟩} completion: completion on file names. If you do not want readline to complete on filenames, you can change the binding of the ^{⟨}\underline{\overline{TAB}}^{⟩} key with rl_bind_key ().

     int rl_bind_key (int key, int (*function)());

rl_bind_key () takes 2 arguments; key is the character that you want to bind, and function is the address of the function to run when key is pressed. Binding ^{⟨}\underline{\overline{TAB}}^{⟩} to rl_ insert () makes ^{⟨}\underline{\overline{TAB}}^{⟩} just insert itself.

rl_bind_key () returns non-zero if key is not a valid ASCII character code (between 0 and 255).

     rl_bind_key ('\t', rl_insert);

This code should be executed once at the start of your program; you might write a function called initialize_readline () which performs this and other desired initializations, such as installing custom completers, etc.

1.2 Custom Functions

Readline provides a great many functions for manipulating the text of the line. But it isn’t possible to anticipate the needs of all programs. This section describes the various functions and variables defined in within the Readline library which allow a user program to add customized functionality to Readline.

1.2.1 The Function Type

For the sake of readabilty, we declare a new type of object, called Function. A Function is a C language function which returns an int. The type declaration for Function is:

typedef int Function ();

The reason for declaring this new type is to make it easier to write code describing pointers to C functions. Let us say we had a variable called func which was a pointer to a function. Instead of the classic C declaration

   int (*)()func;

we have

   Function *func;

1.2.2 Naming a Function

The user can dynamically change the bindings of keys while using Readline. This is done by representing the function with a descriptive name. The user is able to type the descriptive name when referring to the function. Thus, in an init file, one might find

     Meta-Rubout:  backward-kill-word

This binds the keystroke ^{⟨}\underline{\overline{Meta-Rubout}}^{⟩} to the function descriptively named backwardkill-word. You, as the programmer, should bind the functions you write to descriptive names as well. Readline provides a function for doing that:

Function rl add defun (char *name, Function *function, int key)

Add name to the list of named functions. Make function be the function that gets called. If key is not -1, then bind it to function using rl_bind_key ().

Using this function alone is sufficient for most applications. It is the recommended way to add a few functions to the default functions that Readline has built in already. If you need to do more or different things than adding a function to Readline, you may need to use the underlying functions described below.

1.2.3 Selecting a Keymap

Key bindings take place on a keymap. The keymap is the association between the keys that the user types and the functions that get run. You can make your own keymaps, copy existing keymaps, and tell Readline which keymap to use.

Function Keymap rl make bare keymap ()

Returns a new, empty keymap. The space for the keymap is allocated with malloc (); you should free () it when you are done.

Function Keymap rl copy keymap (Keymap map)

Return a new keymap which is a copy of map.

Function Keymap rl make keymap ()

Return a new keymap with the printing characters bound to rl insert, the lowercase Meta characters bound to run their equivalents, and the Meta digits bound to produce numeric arguments.

1.2.4 Binding Keys

You associate keys with functions through the keymap. Here are functions for doing that.

Function int rl bind key (int key, Function *function)

Binds key to function in the currently selected keymap. Returns non-zero in the case of an invalid key.

Function int rl bind key in map (int key, Function *function, Keymap

map) Bind key to function in map. Returns non-zero in the case of an invalid key.

Function int rl unbind key (int key)

Make key do nothing in the currently selected keymap. Returns non-zero in case of error.

Function int rl unbind key in map (int key, Keymap map)

Make key be bound to the null function in map. Returns non-zero in case of error.

Function rl generic bind (int type, char *keyseq, char *data, Keymap

map) Bind the key sequence represented by the string keyseq to the arbitrary pointer data. type says what kind of data is pointed to by data; right now this can be a function (ISFUNC), a macro (ISMACR), or a keymap (ISKMAP). This makes new keymaps as necessary. The initial place to do bindings is in map.

1.2.5 Writing a New Function

In order to write new functions for Readline, you need to know the calling conventions for keyboard invoked functions, and the names of the variables that describe the current state of the line gathered so far.

Variable char *rl line buffer This is the line gathered so far. You are welcome to modify the contents of this, but see Undoing, below.

Variable int rl point The offset of the current cursor position in rl line buffer.

Variable int rl end The number of characters present in rl_line_buffer. When rl_point is at the end of the line, then rl_point and rl_end are equal.

The calling sequence for a command foo looks like

     foo (int count, int key)

where count is the numeric argument (or 1 if defaulted) and key is the key that invoked this function.

It is completely up to the function as to what should be done with the numeric argument; some functions use it as a repeat count, other functions as a flag, and some choose to ignore it. In general, if a function uses the numeric argument as a repeat count, it should be able to do something useful with a negative argument as well as a positive argument. At the very least, it should be aware that it can be passed a negative argument.

1.2.6 Allowing Undoing

Supporting the undo command is a painless thing to do, and makes your functions much more useful to the end user. It is certainly easy to try something if you know you can undo it. I could use an undo function for the stock market.

If your function simply inserts text once, or deletes text once, and it calls rl_insert_ text () or rl_delete_text () to do it, then undoing is already done for you automatically, and you can safely skip this section.

If you do multiple insertions or multiple deletions, or any combination of these operations, you should group them together into one operation. This can be done with rl_begin_undo_

group () and rl_end_undo_group ().

Function rl begin undo group ()

Begins saving undo information in a group construct. The undo information

     usually comes from calls to rl_insert_text () and rl_delete_text (), but

they could be direct calls to rl_add_undo ().

Function rl end undo group ()

Closes the current undo group started with rl_begin_undo_group (). There should be exactly one call to rl_end_undo_group () for every call to rl_begin_

     undo_group ().

Finally, if you neither insert nor delete text, but directly modify the existing text (e.g. change its case), you call rl_modifying () once, just before you modify the text. You must supply the indices of the text range that you are going to modify.

Function rl modifying (int start, int end)

Tell Readline to save the text between start and end as a single undo unit. It is assumed that subsequent to this call you will modify that range of text in some way.

1.2.7 An Example

Here is a function which changes lowercase characters to the uppercase equivalents, and uppercase characters to the lowercase equivalents. If this function was bound to ‘M-c’, then typing ‘M-c’ would change the case of the character under point. Typing ‘10 M-c’ would change the case of the following 10 characters, leaving the cursor on the last character changed.

     /* Invert the case of the COUNT following characters. */
     invert_case_line (count, key)
          int count, key;
     {
       register int start, end;
       start = rl_point;
       if (count < 0)
         {
           direction = -1;
           count = -count;
         }
       else
         direction = 1;

/* Find the end of the range to modify. */
end = start + (count * direction);
/* Force it to be within range. */
if (end > rl_end)
  end = rl_end;
else if (end < 0)
  end = -1;
if (start > end)
  {
    int temp = start;
    start = end;
    end = temp;
  }
if (start == end)
  return;
/* Tell readline that we are modifying the line, so save the undo
   information. */
rl_modifying (start, end);
for (; start != end; start += direction)
  {
    if (uppercase_p (rl_line_buffer[start]))
      rl_line_buffer[start] = to_lower (rl_line_buffer[start]);
    else if (lowercase_p (rl_line_buffer[start]))
      rl_line_buffer[start] = to_upper (rl_line_buffer[start]);
  }
/* Move point to on top of the last character changed. */
rl_point = end - direction;
}

1.3 Custom Completers

Typically, a program that reads commands from the user has a way of disambiguating commands and data. If your program is one of these, then it can provide completion for either commands, or data, or both commands and data. The following sections describe how your program and Readline cooperate to provide this service to end users.

1.3.1 How Completing Works

In order to complete some text, the full list of possible completions must be available. That is to say, it is not possible to accurately expand a partial word without knowing what all of the possible words that make sense in that context are. The GNU Readline library provides the user interface to completion, and additionally, two of the most common completion functions; filename and username. For completing other types of text, you must write your own completion function. This section describes exactly what those functions must do, and provides an example function.

There are three major functions used to perform completion:

1. The user-interface function rl_complete (). This function is called interactively with the same calling conventions as other functions in readline intended for interactive use; i.e. count, and invoking-key. It isolates the word to be completed and calls completion_matches () to generate a list of possible completions. It then either lists the possible completions or actually performs the completion, depending on which behaviour is desired.

2. The internal function completion_matches () uses your generator function to generate the list of possible matches, and then returns the array of these matches. You should place the address of your generator function in rl_completion_entry_function.

3. The generator function is called repeatedly from completion_matches (), returning a string each time. The arguments to the generator function are text and state. text is the partial word to be completed. state is zero the first time the function is called, and a positive non-zero integer for each subsequent call. When the generator function returns (char *)NULL this signals completion_matches () that there are no more possibilities left.

Function rl complete (int ignore, int invoking key)

Complete the word at or before point. You have supplied the function that does the initial simple matching selection algorithm (see completion_matches ()). The default is to do filename completion.

Note that rl_complete () has the identical calling conventions as any other keyinvokable function; this is because by default it is bound to the ‘TAB’ key.

Variable Function *rl completion entry function This is a pointer to the generator function for completion_matches (). If

the value of rl_completion_entry_function is (Function *)NULL then the

default filename generator function is used, namely filename_entry_function

().

1.3.2 Completion Functions

Here is the complete list of callable completion functions present in Readline.

Function rl complete internal (int what to do)

Complete the word at or before point. what to do says what to do with the completion. A value of ‘?’ means list the possible completions. ‘TAB’ means do standard completion. ‘*’ means insert all of the possible completions.

Function rl complete (int ignore, int invoking key)

Complete the word at or before point. You have supplied the function that does the initial simple matching selection algorithm (see completion_matches ()). The default is to do filename completion. This just calls rl_complete_ internal () with an argument of ‘TAB’.

Function rl possible completions ()

List the possible completions. See description of rl_complete (). This just calls rl_complete_internal () with an argument of ‘?’.

Function char **completion matches (char *text, char

*(*entry function) ()) Returns an array of (char *) which is a list of completions for text. If there are no completions, returns (char **)NULL. The first entry in the returned array is the substitution for text. The remaining entries are the possible completions. The array is terminated with a NULL pointer.

entry function is a function of two args, and returns a (char *). The first argument is text. The second is a state argument; it is zero on the first call, and non-zero on subsequent calls. It returns a NULL pointer to the caller when there are no more matches.

Function char *filename completion function (char *text, int state)

A generator function for filename completion in the general case. Note that completion in the Bash shell is a little different because of all the pathnames that must be followed when looking up the completion for a command.

Function char *username completion function (char *text, int state)

A completion generator for usernames. text contains a partial username preceded by a random character (usually ‘~’).

1.3.3 Completion Variables

Variable Function *rl completion entry function A pointer to the generator function for completion_matches (). NULL means to use filename_entry_function (), the default filename completer.

Variable Function *rl attempted completion function A pointer to an alternative function to create matches. The function is called with text, start, and end. start and end are indices in rl_line_buffer saying what the boundaries of text are. If this function exists and returns NULL

then rl_complete () will call the value of rl_completion_entry_function

to generate matches, otherwise the array of strings returned will be used.

Variable int rl completion query items Up to this many items will be displayed in response to a possible-completions call. After that, we ask the user if she is sure she wants to see them all. The default value is 100.

Variable char *rl basic word break characters The basic list of characters that signal a break between words for the completer routine. The contents of this variable is what breaks words in the Bash shell, i.e. " \t\n\"\\’‘@$><=;|&{(".

Variable char *rl completer word break characters The list of characters that signal a break between words for rl_complete_

internal (). The default list is the contents of rl_basic_word_break_
characters.

Variable char *rl special prefixes The list of characters that are word break characters, but should be left in text when it is passed to the completion function. Programs can use this to help determine what kind of completing to do.

Variable int rl ignore completion duplicates If non-zero, then disallow duplicates in the matches. Default is 1.

Variable int rl filename completion desired Non-zero means that the results of the matches are to be treated as filenames. This is always zero on entry, and can only be changed within a completion entry generator function.

Variable Function *rl ignore some completions function This function, if defined, is called by the completer when real filename completion is done, after all the matching names have been generated. It is passed a NULL terminated array of (char *) known as matches in the code. The 1st element (matches[0]) is the maximal substring that is common to all matches. This function can re-arrange the list of matches as required, but each deleted element of the array must be free()’d.

1.3.4 A Short Completion Example

Here is a small application demonstrating the use of the GNU Readline library. It is called fileman, and the source code resides in ‘readline/examples/fileman.c’. This sample application provides completion of command names, line editing features, and access to the history list.

/* fileman.c -- A tiny application which demonstrates how to use the
   GNU Readline library.  This application interactively allows users
   to manipulate files and their modes. */
#include <stdio.h>
#include <readline/readline.h>
#include <readline/history.h>
#include <sys/types.h>
#include <sys/file.h>
#include <sys/stat.h>
#include <sys/errno.h>
/* The names of functions that actually do the manipulation. */
int com_list (), com_view (), com_rename (), com_stat (), com_pwd ();
int com_delete (), com_help (), com_cd (), com_quit ();
/* A structure which contains information on the commands this program
   can understand. */
typedef struct {
  char *name;                   /* User printable name of the function. */
  Function *func;               /* Function to call to do the job. */
  char *doc;                    /* Documentation for this function.  */
} COMMAND;
COMMAND commands[] = {
  { "cd", com_cd, "Change to directory DIR" },
  { "delete", com_delete, "Delete FILE" },
  { "help", com_help, "Display this text" },
  { "?", com_help, "Synonym for `help'" },
  { "list", com_list, "List files in DIR" },
  { "ls", com_list, "Synonym for `list'" },
  { "pwd", com_pwd, "Print the current working directory" },
  { "quit", com_quit, "Quit using Fileman" },
  { "rename", com_rename, "Rename FILE to NEWNAME" },
  { "stat", com_stat, "Print out statistics on FILE" },
  { "view", com_view, "View the contents of FILE" },
  { (char *)NULL, (Function *)NULL, (char *)NULL }
};
/* The name of this program, as taken from argv[0]. */
char *progname;
/* When non-zero, this global means the user is done using this program. */
int done = 0;

main (argc, argv)
   int argc;
   char **argv;
{
progname = argv[0];
initialize_readline ();       /* Bind our completer. */
/* Loop reading and executing lines until the user quits. */
while (!done)
  {
    char *line;
    line = readline ("FileMan: ");
    if (!line)
      {
        done = 1;             /* Encountered EOF at top level. */
      }
    else
      {
        /* Remove leading and trailing whitespace from the line.
           Then, if there is anything left, add it to the history list
           and execute it. */
        stripwhite (line);
        if (*line)
          {
            add_history (line);
            execute_line (line);
          }
      }
    if (line)
      free (line);
  }
exit (0);
}
/* Execute a command line. */
execute_line (line)
   char *line;
{
register int i;
COMMAND *find_command (), *command;
char *word;

/* Isolate the command word. */
i = 0;
while (line[i] && !whitespace (line[i]))
  i++;
word = line;
if (line[i])
  line[i++] = '\0';
command = find_command (word);
if (!command)
  {
    fprintf (stderr, "%s: No such command for FileMan.\n", word);
    return;
  }
/* Get argument to command, if any. */
while (whitespace (line[i]))
  i++;
word = line + i;
/* Call the function. */
(*(command->func)) (word);
}
/* Look up NAME as the name of a command, and return a pointer to that
 command.  Return a NULL pointer if NAME isn't a command name. */
COMMAND *
find_command (name)
   char *name;
{
register int i;
for (i = 0; commands[i].name; i++)
  if (strcmp (name, commands[i].name) == 0)
    return (&commands[i]);
return ((COMMAND *)NULL);
}
/* Strip whitespace from the start and end of STRING. */
stripwhite (string)
   char *string;
{
register int i = 0;

while (whitespace (string[i]))
  i++;
if (i)
  strcpy (string, string + i);
i = strlen (string) - 1;
while (i > 0 && whitespace (string[i]))
  i--;
string[++i] = '\0';
}

/* **************************************************************** */
/*                                                                  */
/*                  Interface to Readline Completion                */
/*                                                                  */
/* **************************************************************** */
/* Tell the GNU Readline library how to complete.  We want to try to complete
   on command names if this is the first word in the line, or on filenames
   if not. */
initialize_readline ()
{
  char **fileman_completion ();
  /* Allow conditional parsing of the ~/.inputrc file. */
  rl_readline_name = "FileMan";
  /* Tell the completer that we want a crack first. */
  rl_attempted_completion_function = (Function *)fileman_completion;
}
/* Attempt to complete on the contents of TEXT.  START and END show the
   region of TEXT that contains the word to complete.  We can use the
   entire line in case we want to do some simple parsing.  Return the
   array of matches, or NULL if there aren't any. */
char **
fileman_completion (text, start, end)
     char *text;
     int start, end;
{
  char **matches;
  char *command_generator ();
  matches = (char **)NULL;
  /* If this word is at the start of the line, then it is a command
     to complete.  Otherwise it is the name of a file in the current
     directory. */
  if (start == 0)
    matches = completion_matches (text, command_generator);
  return (matches);
}
/* Generator function for command completion.  STATE lets us know whether
   to start from scratch; without any state (i.e. STATE == 0), then we
   start at the top of the list. */
char *

command_generator (text, state)
   char *text;
   int state;
{
static int list_index, len;
char *name;
/* If this is a new word to complete, initialize now.  This includes
   saving the length of TEXT for efficiency, and initializing the index
   variable to 0. */
if (!state)
  {
    list_index = 0;
    len = strlen (text);
  }
/* Return the next name which partially matches from the command list. */
while (name = commands[list_index].name)
  {
    list_index++;
    if (strncmp (name, text, len) == 0)
      return (name);
  }
/* If no names matched, then return NULL. */
return ((char *)NULL);
}

/* **************************************************************** */
/*                                                                  */
/*                       FileMan Commands                           */
/*                                                                  */
/* **************************************************************** */
/* String to pass to system ().  This is for the LIST, VIEW and RENAME
   commands. */
static char syscom[1024];
/* List the file(s) named in arg. */
com_list (arg)
     char *arg;
{
  if (!arg)
    arg = "*";
  sprintf (syscom, "ls -FClg %s", arg);
  system (syscom);
}
com_view (arg)
     char *arg;
{
  if (!valid_argument ("view", arg))
    return;
  sprintf (syscom, "cat %s | more", arg);
  system (syscom);
}
com_rename (arg)
     char *arg;
{
  too_dangerous ("rename");
}
com_stat (arg)
     char *arg;
{
  struct stat finfo;
  if (!valid_argument ("stat", arg))
    return;
  if (stat (arg, &finfo) == -1)
    {

    perror (arg);
    return;
  }
printf ("Statistics for `%s':\n", arg);
printf ("%s has %d link%s, and is %d bytes in length.\n", arg,
        finfo.st_nlink, (finfo.st_nlink == 1) ? "" : "s",  finfo.st_size);
printf ("      Created on: %s", ctime (&finfo.st_ctime));
printf ("  Last access at: %s", ctime (&finfo.st_atime));
printf ("Last modified at: %s", ctime (&finfo.st_mtime));
}
com_delete (arg)
   char *arg;
{
too_dangerous ("delete");
}
/* Print out help for ARG, or for all of the commands if ARG is
 not present. */
com_help (arg)
   char *arg;
{
register int i;
int printed = 0;
for (i = 0; commands[i].name; i++)
  {
    if (!*arg || (strcmp (arg, commands[i].name) == 0))
      {
        printf ("%s\t\t%s.\n", commands[i].name, commands[i].doc);
        printed++;
      }
  }
if (!printed)
  {
    printf ("No commands match `%s'.  Possibilties are:\n", arg);
    for (i = 0; commands[i].name; i++)
      {
        /* Print in six columns. */
        if (printed == 6)
          {
            printed = 0;
            printf ("\n");
          }

          printf ("%s\t", commands[i].name);
          printed++;
        }
      if (printed)
        printf ("\n");
    }
}
/* Change to the directory ARG. */
com_cd (arg)
     char *arg;
{
  if (chdir (arg) == -1)
    perror (arg);
  com_pwd ("");
}
/* Print out the current working directory. */
com_pwd (ignore)
     char *ignore;
{
  char dir[1024];
  (void) getwd (dir);
  printf ("Current directory is %s\n", dir);
}
/* The user wishes to quit using this program.  Just set DONE non-zero. */
com_quit (arg)
     char *arg;
{
  done = 1;
}
/* Function which tells you that you can't do this. */
too_dangerous (caller)
     char *caller;
{
  fprintf (stderr,
           "%s: Too dangerous for me to distribute.  Write it yourself.\n",
           caller);
}
/* Return non-zero if ARG is a valid argument for CALLER, else print

   an error message and return zero. */
int
valid_argument (caller, arg)
     char *caller, *arg;
{
  if (!arg || !*arg)
    {
      fprintf (stderr, "%s: Argument required.\n", caller);
      return (0);
    }
  return (1);
}

28 GNU Readline Library Concept Index 29

Concept Index

I

interaction, readline . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

R

readline, function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

V

vi style command editing . . . . . . . . . . . . . . . . . . . . . . . 8

30 GNU Readline Library Function and Variable Index 31

Function and Variable Index

A

abort (⟨\underline{\overline{C-g}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

accept-line (Newline, Return). . . . . . . . . . . . . . . . 5

B

backward-char (⟨\underline{\overline{C-b}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . 5

backward-delete-char (Rubout) . . . . . . . . . . . . . . . 6

backward-kill-line () . . . . . . . . . . . . . . . . . . . . . . . . 7 backward-kill-word (⟨\underline{\overline{M-DEL}}⟩) . . . . . . . . . . . . . . . . 7 backward-word (⟨\underline{\overline{M-b}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . 5

beginning-of-history (⟨\underline{\overline{M-<}}⟩) . . . . . . . . . . . . . . . . . 5

beginning-of-line (⟨\underline{\overline{C-a}}⟩) . . . . . . . . . . . . . . . . . . . . . 5

C

capitalize-word (⟨\underline{\overline{M-c}}⟩) . . . . . . . . . . . . . . . . . . . . . . . 6 char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 16, 17

char **completion matches . . . . . . . . . . . . . . . . . . . 16
char *filename completion function . . . . . . . . . 16
char *username completion function . . . . . . . . . 16

clear-screen (⟨\underline{\overline{C-l}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 complete (TAB) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

D

delete-char (⟨\underline{\overline{C-d}}⟩). . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 digit-argument (⟨\underline{\overline{M-0}}⟩, ⟨\underline{\overline{M-1}}⟩, ... ⟨\underline{\overline{M--}}⟩) . . . . . . 7 downcase-word (⟨\underline{\overline{M-l}}⟩). . . . . . . . . . . . . . . . . . . . . . . . . . 6

E

editing-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 end-of-history (⟨\underline{\overline{M->}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . 5 end-of-line (⟨\underline{\overline{C-e}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

forward-char (⟨\underline{\overline{C-f}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

forward-search-history (⟨\underline{\overline{C-s}}⟩) . . . . . . . . . . . . . . . 5

forward-word (⟨\underline{\overline{M-f}}⟩). . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 16, 17

H

horizontal-scroll-mode . . . . . . . . . . . . . . . . . . . . . . . . . . 4

I

int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 16, 17 int rl bind key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 int rl bind key in map . . . . . . . . . . . . . . . . . . . . . . . 11 int rl unbind key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 int rl unbind key in map . . . . . . . . . . . . . . . . . . . . . 12

K

Keymap rl copy keymap . . . . . . . . . . . . . . . . . . . . . . . . 11

                                                Keymap rl make bare keymap. . . . . . . . . . . . . . . . . . . 11

Keymap rl make keymap . . . . . . . . . . . . . . . . . . . . . . . . 11 kill-line (⟨\underline{\overline{C-k}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 kill-word (⟨\underline{\overline{M-d}}⟩). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

M

mark-modified-lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

N

next-history (⟨\underline{\overline{C-n}}⟩). . . . . . . . . . . . . . . . . . . . . . . . . . . 5

P

                                                possible-completions (M-?). . . . . . . . . . . . . . . . . . . 8

prefer-visible-bell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 prefix-meta (ESC) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 previous-history (⟨\underline{\overline{C-p}}⟩) . . . . . . . . . . . . . . . . . . . . . . 5

Q

quoted-insert (⟨\underline{\overline{C-q}}⟩, ⟨\underline{\overline{C-v}}⟩) . . . . . . . . . . . . . . . . . . . 6

R

re-read-init-file (⟨\underline{\overline{C-x}}⟩ ⟨\underline{\overline{C-r}}⟩). . . . . . . . . . . . . . . . 8 readline () . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

                                                reverse-search-history (⟨\underline{\overline{C-r}}⟩) . . . . . . . . . . . . . . . 5

revert-line (⟨\underline{\overline{M-r}}⟩). . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 rl add defun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 rl_begin_undo_group . . . . . . . . . . . . . . . . . . . . . . . . . 13 rl bind key () . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 rl complete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 rl complete internal . . . . . . . . . . . . . . . . . . . . . . . . . 15 rl_end_undo_group. . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 rl generic bind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 rl_modifying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 rl possible completions . . . . . . . . . . . . . . . . . . . . . 16

S

                                                self-insert (a, b, A, 1, !, ...) . . . . . . . . . . . . . . 6

T

tab-insert (⟨\underline{\overline{M-TAB}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . 6 transpose-chars (⟨\underline{\overline{C-t}}⟩) . . . . . . . . . . . . . . . . . . . . . . . 6 transpose-words (⟨\underline{\overline{M-t}}⟩) . . . . . . . . . . . . . . . . . . . . . . . 6

32 GNU Readline Library

U

undo (⟨\underline{\overline{C-}} ⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 universal-argument () . . . . . . . . . . . . . . . . . . . . . . . . 7 unix-line-discard (⟨\underline{\overline{C-u}}⟩) . . . . . . . . . . . . . . . . . . . . . 7 unix-word-rubout (⟨\underline{\overline{C-w}}⟩). . . . . . . . . . . . . . . . . . . . . . 7 upcase-word (⟨\underline{\overline{M-u}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Y

yank (⟨\underline{\overline{C-y}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 yank-pop (⟨\underline{\overline{M-y}}⟩) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

Table of Contents

Appendix A Command Line Editing . . . . . . . . . . 1

A.1 Introduction to Line Editing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 A.2 Readline Interaction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

A.2.1 Readline Bare Essentials . . . . . . . . . . . . . . . . . . . . . . . . 1 A.2.2 Readline Movement Commands . . . . . . . . . . . . . . . . . 2 A.2.3 Readline Killing Commands. . . . . . . . . . . . . . . . . . . . . 2 A.2.4 Readline Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

A.3 Readline Init File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

A.3.1 Readline Init Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 A.3.1.1 Commands For Moving . . . . . . . . . . . . . . . . 5 A.3.1.2 Commands For Manipulating The History

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 A.3.1.3 Commands For Changing Text . . . . . . . . . 6 A.3.1.4 Killing And Yanking . . . . . . . . . . . . . . . . . . . 7 A.3.1.5 Specifying Numeric Arguments . . . . . . . . . 7 A.3.1.6 Letting Readline Type For You . . . . . . . . . 7 A.3.1.7 Some Miscellaneous Commands. . . . . . . . . 8

A.3.2 Readline vi Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

1 Programming with GNU Readline . . . . . . . . . . 9

1.1 Default Behaviour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 1.2 Custom Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

1.2.1 The Function Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.2.2 Naming a Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.2.3 Selecting a Keymap. . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.2.4 Binding Keys. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.2.5 Writing a New Function. . . . . . . . . . . . . . . . . . . . . . . . 12 1.2.6 Allowing Undoing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.2.7 An Example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

1.3 Custom Completers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 1.3.1 How Completing Works . . . . . . . . . . . . . . . . . . . . . . . . 14 1.3.2 Completion Functions . . . . . . . . . . . . . . . . . . . . . . . . . . 15 1.3.3 Completion Variables . . . . . . . . . . . . . . . . . . . . . . . . . . 16 1.3.4 A Short Completion Example. . . . . . . . . . . . . . . . . . . 17

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

Function and Variable Index . . . . . . . . . . . . . . . . . . 31

ii GNU Readline Library